OpenStudy (anonymous):

Find the composition f(g) of the functions F(x)= 4x-2/x^2 , g(x)=3x+1

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

I think that in your f(x) function, wherever you see x, substitute the g(x) function which is 3x+1

OpenStudy (anonymous):

So what should I do?

OpenStudy (anonymous):

@ybarrap can you please help

OpenStudy (ybarrap):

So, g(x)=3x+1 Using this equation and plug in as the x value for F(x): $$ F(x)= 4x-\cfrac{2}{x^2}\\ $$ So you will have $$ F(g(x))=F(3x+1)=? $$ That's it!

OpenStudy (nikato):

So pluggin it would give u f(g)= 4(3x+1)-2/(3x+1)^2

OpenStudy (nikato):

And you can simplify that if you need to